OpenStudy (anonymous):

Find the equation of the ellipse with the following properties. The ellipse with foci at (0, 6) and (0, -6); y-intercepts (0, 8) and (0, -8).

OpenStudy (mertsj):

a^2-b^2=c^2 a=8 c=6

OpenStudy (mertsj):

b^2=28 Equation: